OpenStudy (anonymous):

How many solutions does the system have? 4y=x−7 5x−8y=14

OpenStudy (princeharryyy):

1

OpenStudy (anonymous):

thanky ou:)

OpenStudy (princeharryyy):

wlcm.!

OpenStudy (anonymous):

r u sure?

OpenStudy (anonymous):

i have a test

OpenStudy (princeharryyy):

ofcouse

OpenStudy (anonymous):

can you help me with my test please?

OpenStudy (princeharryyy):

ohk..

OpenStudy (anonymous):

x+5y =0 15y =−3x

OpenStudy (anonymous):

x+5y=0 15y=−3x

OpenStudy (princeharryyy):

0

OpenStudy (princeharryyy):

no solution

OpenStudy (anonymous):

River C is 300 miles longer than River D. If the sum of their lengths is 5,600 miles, what is the length of each river?

OpenStudy (princeharryyy):

d = 2650 c = 2950

OpenStudy (princeharryyy):

be quick @hibba i don't have enough time

OpenStudy (anonymous):

ok sorry River C is 300 miles longer than River D. If the sum of their lengths is 5,600 miles, what is the length of each river?

OpenStudy (anonymous):

@princeharryyy

OpenStudy (anonymous):

River C is 300 miles longer than River D. If the sum of their lengths is 5,600 miles, what is the length of each river? @princeharryyy

OpenStudy (princeharryyy):

d = 2650 c = 2950

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

using addition x+y=8 x−y=2

OpenStudy (anonymous):

@princeharryyy

OpenStudy (anonymous):

Which of these systems of equations would be best solved using the subtraction method? I. 3x−7y+5y=8 −9x+2y=−38 II. 2y=−8x+9x−2 y=3x−6 III. −7x+5x+3y=8 −2x−9y=−38

OpenStudy (anonymous):

@princeharryyy Which of these systems of equations would be best solved using the subtraction method? I. 3x−7y+5y=8 −9x+2y=−38 II. 2y=−8x+9x−2 y=3x−6 III. −7x+5x+3y=8 −2x−9y=−38

OpenStudy (princeharryyy):

III. −7x+5x+3y=8 −2x−9y=−38